# CORE: Comprehensive Post-OCR Parsing and Receipt Understanding Dataset

### v1

@@ Abstract
In the fields of Optical Character Recognition (OCR) and Natural Language Processing (NLP), integrating multilingual capabilities remains a critical challenge, especially when considering languages with complex scripts such as Arabic. This paper introduces the Comprehensive Post-OCR Parsing and Receipt Understanding Dataset (CORU), a novel dataset specifically designed to enhance OCR and information extraction from receipts in multilingual contexts involving Arabic and English. CORU consists of over 20,000 annotated receipts from diverse retail settings, including supermarkets and clothing stores, alongside 30,000 annotated images for OCR that were utilized to recognize each detected line, and 10,000 items annotated for detailed information extraction. These annotations capture essential details such as merchant names, item descriptions, total prices, receipt numbers, and dates. They are structured to support three primary computational tasks: object detection, OCR, and information extraction. We establish the baseline performance for a range of models on CORU to evaluate the effectiveness of traditional methods, like Tesseract OCR, and more advanced neural network-based approaches. These baselines are crucial for processing the complex and noisy document layouts typical of real-world receipts and for advancing the state of automated multilingual document processing. Our datasets are publicly accessible [^1] https://github.com/Update-For-Integrated-Business-AI/CORU.
1 Introduction
Optical Character Recognition (OCR) [13, 7, 11] is the process of converting images of characters into digital texts. With the advent of deep learning and advancements in the field of computer vision, the performance of OCR has seen significant improvements. However, there is still room for enhancement, particularly in tasks that intersect with Natural Language Processing (NLP) and Information Extraction (IE) [13]. One of the most challenging in this domain is post-OCR parsing. This task involves predicting predefined semantic labels from the OCR output, a process that is far from straightforward. Existing datasets often fall short of providing a comprehensive environment for this research problem. For instance, OCR datasets typically lack parsing class labels. On the other hand, the parsing datasets usually contain error-free texts, which do not accurately represent the error-prone outputs typical of OCR processes. In recent years, researchers from both OCR and NLP communities have been working to address this issue by compiling datasets with defined labels. Notable contributions in this area include the Scanned Receipts OCR and Information Extraction (SROIE) dataset [6] introduced by the ICDAR 2019 Post-OCR Challenge and the subsequent CORD dataset [15]. The above datasets provide a comprehensive collection of receipt images, accompanied by two distinct types of annotations aimed at addressing OCR and parsing problems. The first type of annotation provides box-level text annotations for OCR, while the second type is designed for parsing tasks.
In this paper, we introduce a novel dataset named CORU, standing for Comprehensive Post-OCR Parsing and Receipt Understanding. In the CORU dataset, we have collected a diverse array of receipt images from various retail environments. Figure LABEL:fig:example_images presents examples of these annotated receipt images, highlighting the variety of formats and the complexity of text layouts. To our knowledge, this is the first publicly available dataset explicitly designed to facilitate advanced post-OCR parsing and semantic understanding of receipts, integrating both English and Arabic texts. The detailed contributions of CORU are listed below:
  • Key Information Detection: The first component of our dataset involves 20,000 human-annotated receipts. These annotations target the extraction of key information such as merchant names, dates, receipt numbers, items, and total prices.
  • Large-Scale OCR Dataset: CORU includes a comprehensive set of 30,000 images for OCR tasks. This large-scale component is crucial for training and refining OCR systems to handle the specific challenges presented by receipts, such as diverse layouts and degraded text quality.
  • Detailed Item Analysis: The dataset also provides detailed annotations for 10,000 items, focusing on extracting and classifying item-specific information. It involves classifying and extracting item-specific data, which is crucial for applications requiring item analysis and classification.
  • Baseline Performance Metrics: We establish baseline performances using traditional and advanced neural network approaches for each of the three primary tasks—object detection, OCR, and information extraction.
  • OCR Model: We introduce an OCR model designed to address the unique challenges of extracting text from receipts. This model leverages a combination of convolutional neural networks and bidirectional LSTMs, optimized to handle the diverse and noisy backgrounds often found in receipt images.
2 Related Work
While several datasets cater to various OCR tasks [13, 1], there is a notable shortage of receipt datasets, particularly for Arabic receipts. To highlight the unique features and comprehensive nature of our dataset, Table 1 compares it against existing datasets like SROIE, MC-OCR, UIT, and CORD. One of the earliest and most popular datasets in the scanned receipt domain is the ICDAR 2019 Challenge on Scanned Receipts OCR and Key Information Extraction (SROIE) dataset [6]. This dataset marked a significant advancement in the automated analysis of scanned receipts by introducing a standardized collection of 1,000 annotated receipt images. It focuses on three tasks: text localization, OCR, and key information extraction. These tasks are essential for document analysis systems with substantial commercial potential. The challenge emphasized the unique difficulties inherent in receipt OCR, such as poor scan quality and complex layouts. The CORD (Consolidated Receipt Dataset) dataset [15] addresses the challenge of integrating OCR with NLP tasks like semantic parsing by providing a comprehensive resource for post-OCR parsing. It includes thousands of Indonesian receipt images with box-level text annotations for OCR and multi-level semantic labels for parsing. Unlike traditional datasets, CORD bridges the gap between OCR and parsing, enabling the development of robust models that can handle OCR errors. It also introduces line annotations for converting two-dimensional OCR text into a well-ordered sequence, thereby enhancing parsing performance. The MC-OCR (Mobile Captured Receipt Recognition Challenge) dataset [20], featured at RIVF 2021 conference, includes 2,436 images of Vietnamese receipts captured via mobile devices. This dataset supports two tasks: predicting receipt quality and recognizing key information fields. UIT-MLReceipts dataset [12] addresses the need for a more extensive and carefully labeled dataset for extracting receipt information, overcoming the limitations of existing datasets like SROIE and CORD. With a focus on enriching the available data, UIT-MLReceipts has been compiled by sourcing receipts from various establishments such as restaurants, cafes, bookstores, and supermarkets, ensuring diversity in structure, color, font, and format. Additionally, images from social media groups were incorporated to further enhance dataset variability.
@@ 2 Related Work
Table 1: Comparative overview of the CORU dataset against existing datasets such as SROIE, MC-OCR, UIT, and CORD, highlighting differences in the number of images, categories, and supported tasks. OB refers to Object detection and IE refers to Information Extraction.
@@
Dataset Name # Images # categories OB OCR IE Item IE Language SROIE 1,000 4 $X$ English MC-OCR 2,436 4 $X$ $X$ English & Vietnamese UIT 2,147 4 $X$ English & Vietnamese CORD 1,000 8 $X$ English CORU (ours) 20,000 5 English & Arabic
3 Dataset and Annotations
In this section, we detail the creation of CORU. Our methodology integrates rigorous data collection, annotation, and quality control processes, ensuring that CORU is both diverse and reflective of real-world scenarios. Below we outline the systematic steps in the development of the dataset.
Data Collection
The initial phase involved gathering a large and diverse array of receipts through the DISCO application [^2] https://discoapp.ai/, a platform that facilitates the uploading of receipt images by registered users. We collected over 100k receipts from various commercial settings such as restaurants, supermarkets, and retail stores, ensuring a broad representation of industries, products, and transaction types to mimic the complexity found in everyday transaction documents.
Annotation Guidelines and Team Setup
To ensure consistent and accurate annotations, we developed detailed annotation guidelines. These guidelines defined key elements for annotation, including merchant names, items, total prices, and dates. We assembled a team of expert annotators trained in these guidelines to maintain uniformity throughout the annotation process.
Receipt Selection and Quality Control
Our team carefully reviewed each receipt against specific quality criteria such as text clarity, completeness of the image, and relevance (ensuring documents were receipts and not other types of bills).
Object Detection Annotation
Using the MakeSense tool [^3] https://www.makesense.ai/, annotators constructed bounding boxes around each identified object on the receipts, subsequently saving these annotations in both YOLO and COCO formats. A secondary review team then examined the annotations to rectify any inaccuracies, ensuring the integrity of the data.
OCR Annotation
As part of the OCR task, the dataset includes images annotated specifically for recognizing textual information such as dates, merchant names, and item details. Figure LABEL:fig:ocr_images showcases examples of OCR images, illustrating the challenges and complexity of text recognition. Each identified object or line was passed to a specialized OCR annotation team, which annotated each segment character by character. Due to the challenges posed by bi-directional text in Arabic and left-to-right English, we developed a custom system to annotate each word from left to right, recording the annotations in a structured list that maintains the positional integrity of the words as they appear on the receipt.
Item-Specific Annotation
Detailed annotations were applied to each receipt element such as item names, classifications, quantities, unit sizes, prices, packaging, and brands. A team of five annotators was tasked with this annotation job, using the established guidelines to ensure a consistent and accurate approach across all receipts.
4 Dataset Analysis
In this section, we delve into the statistical and qualitative analysis of the CORU dataset. The cumulative distribution of classes (Figure LABEL:fig:cdf) reveals a progressive increase in class representation, with a steep rise observed at higher class IDs. This indicates a predominance of item classes within the dataset. The distribution across various item classes is depicted in Figure LABEL:fig:class_distribution. It shows significant variation in the frequency of item types, from ’Soft drinks’ and ’Rice, pasta, and noodles’, which are notably prevalent, to less frequent categories like ’Hair & body care’ and ’Tea & Coffee.’ This diversity is important for developing models that are not only accurate but also versatile in handling a wide range of commercial products. Figure LABEL:fig:ocjects presents the frequency of different object classes such as ’Date,’ ’Item,’ ’Merchant,’ ’Receipt Number,’ and ’Total.’ Notably, ’Item’ class entries are most frequent, underscoring the dataset’s design towards detailed, item-level transaction analysis. This focus is instrumental for applications requiring fine-grained analysis like itemized billing and inventory management. The language composition of the dataset is useful for its application in multilingual OCR tasks. As shown in Figure LABEL:fig:language, Arabic is the predominant language, comprising 53.6% of the dataset, followed by English at 26.2%, and mixed-language content at 20.3%. This multilingual aspect demonstrates the dataset’s applicability in diverse linguistic settings, challenging the models to accurately process and understand content in multiple languages and their combinations.
5 Models
In this section, we present a detailed exploration of various deep learning models that have been evaluated using the CORU dataset. Each model has been selected based on its potential to address specific challenges in OCR, object detection, and information extraction from multilingual receipts. We include a mix of traditional approaches, such as Tesseract OCR [16], and advanced neural network architectures, which leverage recent innovations in deep learning to improve accuracy and efficiency. Our evaluation spans a variety of model types, from convolutional neural networks (CNNs) [14] and Transformers to cutting-edge language models, providing insights into their performance across different computational tasks and conditions outlined in the CORU dataset.
5.1 Object Detection Models
Weakly supervised object localization
refers to the task of identifying object locations in images using only image-level labels rather than precise pixel-level annotations [23, 4]. The model learns to detect objects by classifying image patches and determining their relevance to the overall image label, making it a form of multiple instance learning.
CAM [27] revisits the global average pooling layer and shows how it enables the convolutional neural network (CNN) to have remarkable localization ability despite being trained on image-level labels. The key formula can be represented as follows: $f_{k}=\frac{1}{Z}\sum_{i}\sum_{j}f_{k}(i,j)$ where $f_{k}$ is the $k$-th feature map of size $W\times H$ (width and height), and $Z=W\times H$ is the normalization constant.
HAS [17] introduces the ’Hide-and-Seek’ framework that improves object localization by hiding patches in a training image randomly, and forcing the network to seek other relevant parts when the most discriminative part is hidden.
Adversarial Complementary Learning (ACoL) [25] automatically localizes integral objects of semantic interest with weak supervision. The key idea is to hide the most discriminative part from the model to capture the object’s integral extent. The concept behind this solution involves two classifiers working in tandem, with one classifier’s output guiding the erasure operation on the feature maps fed into the other classifier.
Self-produced Guidance (SPG) [26] masks separately the foreground, the object of interest, from the background to provide the classification networks with spatial correlation information of pixels. The concept involves generating guidance masks based on attention maps and using these masks as auxiliary pixel-level supervision to facilitate the training of classification networks.
Attention-based Dropout Layer (ADL) [5] utilizes the self-attention mechanism to process the feature maps of the model. The key concept involves applying a dropout layer based on the attention map of the input, forcing the model to pay attention to less discriminative parts of the object.
CutMix [22] augmentation strategy, where patches are cut and pasted among training images and the ground truth labels are also mixed proportionally to the area of the patches. The key equation can be represented as follows: CutMix $\text{CutMix Image}=M\odot\text{Image}_{1}+(1-M)\odot\text{Image}_{2}$ where $M$ is a binary mask indicating the cut and paste regions, $\odot$ denotes element-wise multiplication, and Image1 and Image2 are the original and target images, respectively.
DETR [3] the vision transformer model with Improved Denoising Anchor Boxes (DINO) [24] employs a Transformer-based architecture that dynamically updates 4D anchor boxes through its decoder, using deformable attention to leverage multi-scale features effectively. DINO accelerates convergence through a contrastive denoising training method, optimizing detection by managing hard negatives with hyper-parameters $\lambda_{1}$ and $\lambda_{2}$. The mathematical expression for the anchor tracking discrepancy (ATD) is given by: $\text{ATD}(k)=\frac{1}{k}\sum_{\text{topK}}\left\{\|b_{0}-a_{0}\|_{1},\|b_{1}-a_{1}\|_{1},\ldots,\|b_{N-1}-a_{N-1}\|_{1}\right\}$ Additionally, the look-forward-twice strategy updates bounding box predictions iteratively as follows: $\Delta b_{i}=\text{Layer}_{i}(b_{i-1}),\quad b^{\prime}_{i}=\text{Update}(b_{i-1},\Delta b_{i})$ and $b_{i}=\text{Detach}(b^{\prime}_{i}),\quad b_{i}^{\text{pred}}=\text{Update}(b^{\prime}_{i-1},\Delta b_{i})$ This method ensures precise object localization, particularly useful in complex applications such as receipt item detection.
One-Stage Object Detection The advancements in one-stage object detection have been notably driven by the introduction of YOLOv7 [21] and YOLOv8 [10], marking significant milestones in real-time object detection technologies. YOLOv7 improves upon its predecessors by implementing the Focus (Fu) module, which utilizes spatial attention mechanisms to dynamically enhance focus on relevant regions within feature maps, leading to more precise object localization. This is expressed in the equation $F_{k}(i,j)=\sigma(\sum_{m=1}^{C}w_{m}^{k}f_{m}(i,j))$, where each component plays a specific role in enhancing detection precision. The model also incorporates a Path Aggregation Network (PAN) and Efficient Bottleneck ($E-B$) to further optimize performance across various object sizes and improve computational efficiency. On the other hand, YOLOv8 employs CSPDarknet53 as its backbone, which is crucial for multi-scale feature extraction and includes the newly introduced C2f module, enhancing the model’s non-linear representation capabilities. This backbone processes input images at multiple resolutions to detect objects at different scales with high accuracy, and its design incorporates multiple bottleneck blocks that split and concatenate feature maps, optimizing the final convolution layer outputs for accurate detection across diverse object instances. Together, YOLOv7 and YOLOv8 represent a leap forward in the field, combining innovative architectural elements to set new benchmarks for speed and accuracy in object detection.
5.2 OCR
The OCR model combines convolutional neural networks (CNNs) and Long Short-Term Memory (LSTM) units to enhance text extraction and recognition capabilities for optical character recognition (OCR) tasks. We build an OCR Model which contains a convolutional layer designed to perform initial feature extraction. This layer applies multiple filters to the input image, capturing essential textual features such as edges and textural details. The convolution operation can be mathematically represented as: $f(x,y)=\sum_{i=-a}^{a}\sum_{j=-b}^{b}k(i,j)\cdot g(x-i,y-j)$ where $f(x,y)$ denotes the output feature map, $g(x,y)$ the input image, and $k(i,j)$ the convolutional filter. This operation is critical for detecting patterns that form the basis for character recognition. Following the convolutional layer, the extracted features are processed by LSTM units. These units are crucial for managing the sequence data, effectively capturing both short-term and long-term textual dependencies. Using bidirectional LSTMs ensures that the context from all directions is considered, enhancing the model’s ability to decode text sequences accurately.
5.3 LLM models
Llama [18] [^4] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf is a part of the Llama 2 family of Large Language Models (LLMs) developed by Meta. These models are pre-trained and fine-tuned generative text models, specifically optimized for dialogue use cases. They have demonstrated superior performance on numerous benchmarks. The models employ supervised fine-tuning and reinforcement learning with human feedback to align with human preferences for helpfulness and safety. Mistral [8] [^5] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, a model released by Mistral AI, is renowned for its power and efficiency. It surpasses the Llama 2 13B on all benchmarks. The model leverages instruction fine-tuning, where the prompt should be surrounded by [INST] and [/INST] tokens. Mixtral [9] [^6] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, another innovation from Mistral AI, is a trained generative Sparse Mixture of Experts that outperforms the Llama 2 70B model on most benchmarks. The model leverages up to 45B parameters but only uses about 12B during inference, leading to better inference throughput at the cost of more vRAM. Falcon [2] [^7] https://huggingface.co/docs/transformers/main/model_doc/falcon is a class of causal decoder-only models. The largest Falcon checkpoints have been trained on more than 1T tokens of text, with a particular emphasis on the RefinedWeb corpus. Falcon architecture is modern and optimized for inference, with multi-query attention and support for efficient attention variants like FlashAttention. Zephyr [19] [^8] https://huggingface.co/docs/transformers/main/model_doc/zephyr is a state-of-the-art model developed by the Hugging Face. It is designed for a wide range of natural language processing tasks, including text generation, translation, and summarization. The model leverages the Transformer architecture, which allows it to capture long-range dependencies in the input text. The Zephyr model is pre-trained on a large corpus of text and can be fine-tuned for specific tasks.
6 Experimental Results
In this section, we discuss our experimental results using the models described above.
6.1 Object Detection Results
The object detection models we evaluated include Class Activation Mapping (CAM), Hide-and-Seek (HAS), Attention-based Dropout Layer (ADL), Adversarial Complementary Learning (ACOL), Self-produced Guidance (SPG), CutMix, and DETR with Improved Denoising Anchor Boxes (DINO), each employing different architectures like ResNet50, VGG16, InceptionV2, and Swin Transformer. Our evaluation focuses on both classification and localization metrics. Table 2 shows a detailed comparison of object detection performance across various models and backbone configurations. For instance, the DINO model equipped with the Swin 4-scale backbone shows a better average score of 32.2, with notable robustness across different quality thresholds—45.4 at 10%. In contrast, traditional models like CAM and HAS, while effective to a certain degree, have limitations with ResNet50 scoring an average of 6.17 and 7.14, respectively, with performance dropping significantly at higher quality thresholds. Our evaluation of the YOLO models provided significant insights into real-time object detection capabilities within the CORU dataset. In Table 6, YOLOv7 demonstrated the highest performance metrics in terms of precision (P), recall (R), and mean Average Precision (mAP) at IoU=0.50 (mAP50), pointing to robustness in detecting various objects on multilingual receipts with diverse layouts. Although YOLOv8 showed slightly lower performance compared to YOLOv7, it offered improvements in mAP across different IoU thresholds (mAP50-95), suggesting better generalization across diverse object sizes and shapes. YOLOv9 further improved upon these metrics, highlighting the continuous advancements in the YOLO architecture to enhance detection accuracy and efficiency in challenging OCR scenarios.
Table 2: Comparative Analysis of Object Detection Models Across Different Backbone Architectures. This table provides detailed metrics for various object detection models using backbone networks such as ResNet50, VGG16, InceptionV2, and Swin Transformer. Each row represents a different model, evaluating performance at multiple quality thresholds to illustrate their effectiveness in localizing key objects within complex receipt images.

### v2

# ReceiptSense: Beyond Traditional OCR - A Dataset for Receipt Understanding

@@ Abstract.
Multilingual OCR and information extraction from receipts remains challenging, particularly for complex scripts like Arabic. We introduce ReceiptSense, a comprehensive dataset designed for Arabic-English receipt understanding comprising 20,000 annotated receipts from diverse retail settings, 30,000 OCR-annotated images, and 10,000 item-level annotations, and a new Receipt QA subset with 1265 receipt images paired with 40 question-answer pairs each to support LLM evaluation for receipt understanding. The dataset captures merchant names, item descriptions, prices, receipt numbers, and dates to support object detection, OCR, and information extraction tasks. We establish baseline performance using traditional methods (Tesseract OCR) and advanced neural networks, demonstrating the dataset’s effectiveness for processing complex, noisy real-world receipt layouts. Our publicly accessible dataset advances automated multilingual document processing research [^1] https://github.com/Update-For-Integrated-Business-AI/CORU.
Receipt Understanding, Multilingual OCR, Information Extraction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision tasksccs: Computing methodologies Computer visionccs: Computing methodologies Information extraction
1. Introduction
Optical Character Recognition (OCR) (Nguyen et al., 2021a; Hwang et al., 2019; Nguyen et al., 2021b) converts images of characters into digital text. While deep learning has improved OCR performance, challenges remain in post-OCR parsing—predicting semantic labels from noisy, unstructured OCR output, especially for receipts with diverse layouts, fonts, and degraded quality. Existing datasets fall short: standard OCR datasets lack parsing labels, while parsing datasets contain clean text that does not reflect real OCR errors. Recent efforts like SROIE (Huang et al., 2019) and CORD (Park et al., 2019) offer annotated receipts but do not comprehensively address multilingual understanding. We introduce ReceiptSense, the first public dataset for advanced multilingual post-OCR parsing, featuring diverse Arabic-English receipts with detailed OCR, key information, and item-level annotations. Its four components—text-level annotations, key information labels, and item-level data—support OCR enhancement, information extraction, QA and multilingual text analysis. Figure 1 shows examples highlighting the variety of formats and complex layouts.
The primary contributions of the ReceiptSense dataset are as follows: (1) Key Information Detection: 20,000 human-annotated receipts extracting merchant names, dates, receipt numbers, item lists, and totals for key-value information extraction from noisy text. (2) Large-Scale OCR Dataset: 30,000 receipt images with box-level text annotations addressing challenges from diverse layouts, fonts, and noisy appearances. (3) Detailed Item Analysis: 10,000 individual items annotated for item-level analysis in expense management, inventory tracking, and retail analytics. (4) Receipt QA: includes 1,265 receipt images, where each is paired with 40 question-answer pairs covering merchant, date, totals, items, taxes, and payment methods. (5) Baseline Performance Metrics: Comprehensive benchmarks using traditional and modern deep learning approaches across object detection, OCR, and information extraction tasks.
ReceiptSense introduces several novel contributions compared to existing datasets (SROIE, CORD, UIT, MC-OCR). It is significantly larger with 20,000 annotated receipts, 50,600 QA, 30,000 OCR-annotated images, and 10,000 item-level annotations. Unlike existing English-only datasets with clean text, ReceiptSense incorporates multilingual Arabic-English receipts with diverse layouts, fonts, and mixed-language content reflecting global retail scenarios. The dataset includes real-world noise—blurred, rotated, or damaged text—challenging models for robust performance. Crucially, ReceiptSense is the first to provide item-level annotations, enabling fine-grained extraction for expense management and inventory analysis. These characteristics advance multilingual OCR, post-OCR parsing, and practical applications in finance and retail.
Figure 1. Examples of annotated receipt images from the ReceiptSense dataset.
2. Related Work
While several datasets cater to various OCR tasks (Nguyen et al., 2021a; abdallah2024transformers), there is a notable shortage of receipt datasets, particularly for Arabic receipts. To highlight the unique features and comprehensive nature of our dataset, Table 1 compares it against existing datasets like SROIE, MC-OCR, UIT, and CORD.
One of the earliest and most popular datasets in the scanned receipt domain is the ICDAR 2019 Challenge on Scanned Receipts OCR and Key Information Extraction (SROIE) dataset (Huang et al., 2019). This dataset marked a significant advancement in the automated analysis of scanned receipts by introducing a standardized collection of 1,000 annotated receipt images. It focuses on three tasks: text localization, OCR, and key information extraction. These tasks are essential for document analysis systems with substantial commercial potential. The challenge emphasized the unique difficulties inherent in receipt OCR, such as poor scan quality and complex layouts. SROIE has since become a cornerstone for research in receipt document understanding, serving as a reliable benchmark for evaluating the performance of novel models in various OCR and information extraction tasks.
The CORD (Consolidated Receipt Dataset) dataset (Park et al., 2019) addresses the challenge of integrating OCR with NLP tasks like semantic parsing by providing a comprehensive resource for post-OCR parsing. It includes thousands of Indonesian receipt images with box-level text annotations for OCR and multi-level semantic labels for parsing. Unlike traditional datasets, CORD bridges the gap between OCR and parsing, enabling the development of robust models that can handle OCR errors. It also introduces line annotations for converting two-dimensional OCR text into a well-ordered sequence, thereby enhancing parsing performance. CORD’s hierarchical labeling structure allows researchers to investigate both low-level text extraction and high-level semantic interpretation, making it a versatile resource for document understanding tasks.
The MC-OCR (Mobile Captured Receipt Recognition Challenge) dataset (vu2021mc), featured at RIVF 2021 conference, includes 2,436 images of Vietnamese receipts captured via mobile devices. This dataset supports two tasks: predicting receipt quality and recognizing key information fields. Collected by 50 data collectors, these images were annotated in two phases: image quality assessment (IQA) and key information extraction (KIE). IQA involved evaluating text line readability, while KIE required annotators to identify and transcribe key fields. This dataset serves as a benchmark for improving document digitalization and automated financial document processing, with evaluation metrics including RMSE for IQA and CER for KIE. Its mobile-captured nature makes it particularly relevant to real-world applications where varying lighting conditions and image distortions are common challenges.
The UIT-MLReceipts dataset (nguyen2024uit) addresses the need for a more extensive and carefully labeled dataset for extracting receipt information, overcoming the limitations of existing datasets like SROIE and CORD. With a focus on enriching the available data, UIT-MLReceipts has been compiled by sourcing receipts from various establishments such as restaurants, cafes, bookstores, and supermarkets, ensuring diversity in structure, color, font, and format. Additionally, images from social media groups were incorporated to further enhance dataset variability. Following meticulous curation, a total of 2,147 receipt images were obtained and annotated, covering key information like store names, addresses, timestamps, and total costs. These annotations were refined using a Faster R-CNN model trained on the MC-OCR challenge dataset, ensuring accuracy and completeness. The resulting UIT-MLReceipts dataset offers a comprehensive representation of receipt characteristics, languages, status, and image attributes, making it a valuable asset for research in receipt information extraction and Visual Document Understanding (VDU).
Despite the existence of these datasets, there remains a substantial gap in resources tailored for Arabic receipt information extraction. Arabic receipts often present unique challenges such as right-to-left text orientation, diverse numeral systems, and variations in layout structures. Our proposed CORU dataset aims to address these challenges by offering a sizable, high-quality collection of Arabic and English receipts annotated with key information fields. This dataset not only facilitates OCR tasks but also supports advanced question-answering models by including context-aware questions related to receipt content. By bridging this gap, CORU contributes to the broader effort of developing multilingual and multimodal OCR systems capable of handling diverse document structures across languages and regions.
@@ 2. Related Work
Table 1. Comparative overview of the ReceiptSense dataset, highlighting differences in the number of images, categories, and supported tasks. OB refers to Object detection and IE refers to Information Extraction.
@@
Dataset Name # Images # categories OB OCR IE Item IE ReceiptSense Language SROIE 1,000 4 $X$ English MC-OCR 2,436 4 $X$ $X$ English & Vietnamese UIT 2,147 4 $X$ English & Vietnamese CORD 1,000 8 $X$ English ReceiptSense (ours) 20,000 5 English & Arabic
3. Dataset Creation and Analysis
3.1. Data Collection
We developed ReceiptSense to advance multilingual receipt understanding, particularly for Arabic and English texts. Our methodology integrates rigorous data collection, annotation, and quality control processes to ensure diversity and real-world applicability. Ethical Considerations and Data Collection: All receipts were collected with explicit user consent through the DISCO application [^2] https://discoapp.ai/, following strict privacy protocols. We gathered over 100k receipts from restaurants, supermarkets, and retail stores across different geographical regions. To protect privacy, our annotation team applied a four-step PII redaction process: line-by-line review, sensitive data obscuring, verification, and independent cross-checking. Annotation Process: We established detailed guidelines for annotating merchant names, items, prices, and dates, with specialized protocols for bi-directional Arabic-English text. Using MakeSense [^3] https://www.makesense.ai/, annotators created bounding boxes in YOLO and COCO formats. For OCR tasks, we developed a custom system maintaining positional integrity of mixed-language text. Item-specific annotations covered names, classifications, quantities, prices, packaging, and brands, validated through iterative feedback loops with domain experts. For the Receipt QA subset, we used 1,265 real receipt images and manually authored 40 diverse QA pairs per receipt. Questions cover receipt meta-data, item-level details, and transaction summaries. The QA pairs were validated for consistency with the receipt content.
3.2. Dataset Statistical Analysis
ReceiptSense demonstrates significant diversity across multiple dimensions. The language distribution shows Arabic predominance (53.6%), followed by English (26.2%) and mixed-language content (20.3%), reflecting real-world multilingual scenarios in international retail environments. Object class analysis reveals Item entries as most frequent, supporting detailed transaction analysis for applications like itemised billing and inventory management. The item class distribution spans from prevalent categories like ’Soft drinks’ and ’Rice, pasta, and noodles’ to specialised items like ’Hair & body care’, ensuring model versatility across commercial domains. The Receipt QA subset comprises over 50,000 QA pairs (1,265 receipts × 40 questions). The QA coverage includes 30% merchant/payment/date metadata, 50% item-level information, and 20% tax/total/payment details, ensuring comprehensive evaluation of receipt understanding.
4. Models
Visual QA Models
To evaluate receipt-level question answering, we applied state-of-the-art large language models (LLMs) including GPT-4o (Achiam et al., 2023), Llama 3.2 (Touvron et al., 2023a), Phi3 (Abdin et al., 2024), Phi3.5 (Abdin et al., 2024), Llava (Liu et al., 2023), Internvl2 4B (Chen et al., 2024), and Internvl2 8B (Chen et al., 2024). These models were tasked with predicting answers to 40 distinct question types per receipt, covering receipt metadata, item details, VAT, totals, and payment methods.
Object Detection Models:
Weakly supervised models localize objects using only image-level labels, reducing annotation effort while maintaining accuracy (Zhang et al., 2021; Choe et al., 2020). Techniques include Class Activation Mapping (CAM) (Zhou et al., 2016), Hide-and-Seek (HAS) (Singh and Lee, 2017), Adversarial Complementary Learning (ACoL) (Zhang et al., 2018a), Self-produced Guidance (SPG) (Zhang et al., 2018b), Attention-based Dropout Layer (ADL) (Choe and Shim, 2019), and CutMix (Yun et al., 2019), each enhancing feature learning through masking, attention, or augmentation. Advanced architectures like DINO (Zhang et al., 2022) use Transformer-based deformable attention for precise localization, while YOLO models (YOLOv7 (Wang et al., 2023), YOLOv8 (Jocher et al., 2023)) offer real-time detection with multi-scale and attention mechanisms.
OCR Models
Our OCR model combines CNNs and bidirectional LSTMs for text recognition. The convolutional layer performs feature extraction through: $f(x,y)=\sum_{i=-a}^{a}\sum_{j=-b}^{b}k(i,j)\cdot g(x-i,y-j)$, where $f(x,y)$ is the output feature map, $g(x,y)$ the input image, and $k(i,j)$ the convolutional filter. LSTM units capture sequential dependencies for accurate text decoding, with bidirectional processing ensuring comprehensive context understanding.
Large Language Models
We evaluate several LLMs for information extraction tasks. Llama (Touvron et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024), Falcon (Almazrouei et al., 2023) and Zephyr (Tunstall et al., 2023).
Table 2. Comparative Analysis of Object Detection Models Across Different Backbone Architectures.